OpenStudy (anonymous):

when solving for this types of problems does it matter the order you put the numbers? do you always start with the number to the right of the circle is there a rule ? Find the length of segment AC. Round your answer to two decimal places. Segment AB extended is the diameter. Segment CD=21 Segment AB=11

OpenStudy (anonymous):

21^2 + ? = 11 ^2 11^2 + ? = 21^2 that sort of thing

OpenStudy (anonymous):

i dont think it would really matter there but when il subtract to get my answer i think it matters